# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ Abstract
We consider maximal operators acting on vector valued functions, that is functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts. For the reader’s convenience, we include an appendix explaining how to adapt the extrapolation for real convex body valued functions of Bownik and Cruz-Uribe to the setting of complex convex body valued functions that we treat.
@@ 1 Introduction
The classical Fefferman–Stein vector valued inequalities, first proved by C. Fefferman and E. Stein [Fefferman_Stein_1971], state that for all $1<p,q<\infty$ one has
@@
where the supremum ranges over all cubes $Q$ in $\mathbb{R}^{n}$ (with faces parallel to the coordinate hyperplanes), $|E|$ is the Lebesgue measure of a Lebesgue measurable set $E\subseteq\mathbb{R}^{n},$ and $\,\mathrm{dm}(x)$ denotes integration with respect to Lebesgue measure. The constant $C(n,p,q)$ in (1.1) depends only on $n,$ $p$ and $q.$ In [Fefferman_Stein_1971] also a weak type version of (1.1) is proved.
@@
for any sequence $\{f_{k}\}^{\infty}_{k=1}.$ Such inequalities seem to have been studied for the first time systematically by J.-L. Rubio de Francia [Rubio_de_Francia_1985]. For a thorough modern exposition of the methods in [Rubio_de_Francia_1985] we refer to [Tao_notes_5_247A] as well as [Grafakos_Book_1].
The ideas in [Rubio_de_Francia_1985] already hinted at an intimate connection between extrapolation and bounds for vector valued extensions as in (1.2). By an extrapolation problem one understands the following. Given an operator $T$ acting on (suitable) functions on $\mathbb{R},$ we assume that for some $1<p<\infty$ it is already known or given that for all weights (that means, a.e. positive locally integrable functions) $w$ on $\mathbb{R}^{n}$ that belong to some class $C(p),$ one has the estimate $\|Tf\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,w)\|f\|_{\mathrm{L}^{p}(w)}.$ Given this information, find all $1<q<\infty$ and as well as associated classes $C(q)$ of weights on $\mathbb{R}^{n},$ such that for any $w\in C(q)$ one has an estimate of the form $\|Tf\|_{\mathrm{L}^{q}(w)}\leq C(T,n,q,w)\|f\|_{\mathrm{L}^{q}(w)}.$ Rubio de Francia [Rubio_de_Francia_1984] solved completely the extrapolation problem in the case that $C(p)$ coincides with the Muckenhoupt $A_{p}$ class, that is $w\in A_{p}$ if and only if
$[w]_{A_{p}}\coloneq\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}w(x)\,\mathrm{dm}(x)\right)\left(\frac{1}{|Q|}\int_{Q}w(x)^{-1/(p-1)}\,\mathrm{dm}(x)\right)^{p-1}<\infty,$
where the supremum ranges again over all cubes $Q\subseteq\mathbb{R}^{n}.$ In this case, beginning with any fixed $1<p<\infty$ and an extrapolation hypothesis holding for all weights $w\in A_{p},$ the extrapolation problem is solvable for any $1<q<\infty$ and all weights $w\in A_{q}.$ The extrapolation theorem of Rubio de Francia was subsequently further refined by various authors, until a sharp quantitative version of it was proved by J. Duoandikoetxea [Duoandikoetxea_2011]. A very thorough treatment of various forms of extrapolation with extensive historical background can be found in [Book_Extrapolation]. In fact, this method is so powerful, that it naturally yields weighted estimates for vector valued extensions, that is
@@
as explained in detail in [Book_Extrapolation].
Inequalities of the form (1.3) are a major tool for estimating operators arising naturally when decomposing biparameter operators or bicommutators in simpler, localized pieces. Such a decomposition for the so called Journeé operators was established by H. Martikainen [Martikainen_2012] (generalizing an analogous decomposition proved earlier by T. Hytönen [Hytonen_2012] for Calderón–Zygmound operators in the context of the solution of the $A_{2}$ problem). I. Holmes, S. Petermichl and B. Wick [Holmes_Petermichl_Wick_2018] showed that these localized pieces can be estimated in terms of a weighted product BMO space. In the following, we recall the relevant definitions.
@@
where the supremum is taken over all cubes $Q\subseteq\mathbb{R}^{n},$ and we denote $\langle b\rangle_{Q}\coloneq\frac{1}{|Q|}\int_{Q}b(x)\,\mathrm{dm}(x).$ The importance of this space is two-fold. First, it is the dual space to the (real variable) Hardy space $\mathrm{H}^{1}(\mathbb{R}^{n}).$ Second, the norm $\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}$ is the “correct” quantity controlling the boundedness of commutators $[T,b]=[T,M_{b}],$ where $M_{b}$ denotes (pointwise) multiplication by $b$ (called symbol of the commutator) and $T$ is a Calderón–Zygmound operator. This was proved for the Hilbert transform by Z. Nehari [Nehari_1957] and in full generality by R. R. Coifmann, Rochberg and G. Weiss [Coifman_Rochberg_Weiss_1976]. Moreover, the John–Nirenberg inequalities are an important intrinsic property of the space $\mathrm{BMO}(\mathbb{R}^{n}).$
B. Muckenhoupt and R. L. Wheeden [Muckenhoupt_Wheeden_BMO_1976] considered and studied the weighted BMO norm
@@
where the supremum ranges over all cubes $Q\subseteq\mathbb{R}^{n}$ and $\nu$ is a $A_{2}$ weight on $\mathbb{R}^{n}.$ A characterization of the two weighted boundedness of commutators in terms of a weighted BMO norm of the symbol was established by S. Bloom [Bloom_1985] for the Hilbert transform and later for arbitrary Calderón–Zygmund operators by I. Holmes, M. Lacey and B. Wick [Holmes_Lacey_Wick_2016]. In the latter work two weighted versions of (1.4) were introduced and associated John–Nirenberg inequalities were established. These played an important role in the commutator estimates in [Holmes_Lacey_Wick_2016].
The study of biparameter BMO spaces on product spaces $\mathbb{R}^{n}\times\mathbb{R}^{m}$ was initiated by S. Y. A. Chang [Chang_1979] and R. Fefferman [RFefferman_1979]. Here, “biparameter” refers to invariance of the considered function spaces under rescaling each coordinate variable of the domain of definition separately. Works [Chang_1979] and [RFefferman_1979] introduced and investigated the biparameter product BMO space $\text{BMO}(\mathbb{R}\times\mathbb{R})$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{2}$ (considered as the product space $\mathbb{R}\times\mathbb{R}$) such that
@@
The aforementioned works [Chang_1979] and [RFefferman_1979] established in particular that $\text{BMO}(\mathbb{R}\times\mathbb{R})$ is the dual to the biparameter Hardy space $\mathrm{H}^{1}(\mathbb{R}\times\mathbb{R}).$ Moreover, a dyadic version of this product BMO space is the correct space for characterizing the boundedness of bicommutators $[T_{1},[T_{2},b]],$ where $T_{1},T_{2}$ are Haar multipliers, as proved by [Blasco_Pott_2005].
A weighted version of the Chang–Fefferman product BMO space was introduced and studied by Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018] in the context of proving two weight upper bounds for biparameter praproducts. Namely, given a biparameter dyadic grid $\bm{\mathcal{D}}$ in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and a biparameter $\bm{\mathcal{D}}$-dyadic $A_{2}$ weight $w$ on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ [Holmes_Petermichl_Wick_2018] considers the dyadic Bloom type product space $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with the property
@@
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ of nonzero finite measure. We refer to Section 2 for a detailed explanation of the notation and the terminology. In [Holmes_Petermichl_Wick_2018] an $\mathrm{H}^{1}$-BMO duality type result was established in this setting, which played a crucial role in the proofs of the upper bounds there. More recently, a two-weight version of (1.5) was defined in [Kakaroumpas_Soler_2022] and associated John–Nirenberg inequalities were established. These played an important role in [Kakaroumpas_Soler_2022] for characterizing the two weight boundedness of bicommutators with Haar multipliers, extending the aforementioned result of [Blasco_Pott_2005] to the two weight setting.
@@ 1.1 Main results
We refer to Subsection 3.1 for a detailed explanation of the notation in Theorem 1. Theorem 1 yields readily a similar estimate for the so called modified Christ–Goldberg maximal function, as explained in Theorem 7 below.
@@
We deduce Theorem 1 from a general principle for establishing matrix weighted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 5 below. Our method for deducing such bounds is inspired from [Book_Extrapolation]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [Cruz_Uribe_Bownik_Extrapolation], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the recent limited range extrapolation theorem for matrix weights proved in [Kakaroumpas_Nguyen_Vardakis_2024] yields similar bounds as in Theorem 5 that are valid only for a limited range of exponents. Note that the extrapolation for matrix weights and the other methods and techniques from [Cruz_Uribe_Bownik_Extrapolation] (which also belong to the foundations of the work in [Kakaroumpas_Nguyen_Vardakis_2024]) concern the setting of real convex body valued functions. An extension to complex convex body valued functions presents nevertheless no difficulties thanks to previous work in [DKPS2024]. For the reader’s convenience we supply details in the appendix.
@@
Theorem 1 and its counterpart for the modified Christ–Goldberg maximal function, Theorem 7, allow one to complete the proof of matrix weighted bounds for general (not necessarily paraproduct free) Journé operators given in [DKPS2024] for any $1<p<\infty$ (not necessarily $p=2$), as explained in [DKPS2024, Section 8] (thus avoiding the use of extrapolation for biparameter matrix weights as in [Vuorinen_2023]). In this paper we focus on the application of Theorem 1 for setting up the foundations of a theory of two matrix weighted product BMO. Namely, let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We define
$\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\coloneq\sup_{\Omega}\frac{1}{|\Omega|^{1/2}}\bigg{(}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|^{2}\bigg{)}^{1/2},$
@@
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure. This definition is an extension of one of the equivalent definitions for the space of two matrix weighted one-parameter BMO, whose study was initiated in [Isralowitz_Kwon_Pott2017], [Isralowitz_2017] and culminated in [Isralowitz_Pott_Treil_2022]. Moreover, for every sequence $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C}),$ we define
@@
which is the direct biparameter analog of the one parameter two matrix weighted $\mathrm{H}^{1}$ norm from [Isralowitz_2017]. In this context, we prove the following result.
@@
is well-defined and bounded with $\|\ell\|\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$ where the implied constants depend only on $n,m,p,d$ and $[V]_{A_{p},\bm{\mathcal{D}}}.$ Conversely, for every bounded linear functional $\ell$ on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ there is $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with $\ell=\ell_{B}.$
We note that our proof of Theorem 2 trivially works also in the one-parameter setting, thus answering to the positive the question posed in [Isralowitz_2017] about the extension of the one parameter two matrix weighted $\mathrm{H}^{1}$-BMO duality proved there from $p=2$ to arbitrary exponents $1<p<\infty.$
The duality result of Theorem 2 coupled with Theorem 1 yields readily two-matrix weighted bounds for biparameter paraproducts, see Proposition 11 and Proposition 14 below. These yield in turn already some two matrix weighted upper bounds for bicommutators, see Subsection 4.3 below.

### v3

@@ Abstract
We consider maximal operators acting on vector valued functions, that is, functions taking values on $\mathbb{C}^{d},$ that incorporate matrix weights in their definitions. We show vector valued estimates, in the sense of Fefferman–Stein inequalities, for such operators. These are proven using an extrapolation result for convex body valued functions due to Bownik and Cruz-Uribe. Finally, we show an $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality for matrix valued functions and we apply the previous vector valued estimates to show upper bounds for biparameter paraproducts. For the reader’s convenience, we include an appendix explaining how to adapt the extrapolation for real convex body valued functions of Bownik and Cruz-Uribe to the setting of complex convex body valued functions that we treat.
2020 Mathematics Subject Classification: 42B25, 42B30, 42B35.
Keywords: convex set valued functions, maximal operators, matrix weights, Fefferman–Stein vector valued estimates, paraproducts, matrix weighted matrix valued $\mathrm{BMO},$ $\mathrm{H}^{1}$-$\mathrm{BMO}$ duality.
@@ 1 Introduction
The classical Fefferman–Stein vector valued inequalities, first proved by C. Fefferman and E. Stein [17], state that for all $1<p,q<\infty$ one has
@@
where the supremum ranges over all cubes $Q$ in $\mathbb{R}^{n}$ (with faces parallel to the coordinate hyperplanes), $|E|$ is the Lebesgue measure of a Lebesgue measurable set $E\subseteq\mathbb{R}^{n},$ and $\mathrm{dm}(x)$ denotes integration with respect to Lebesgue measure. The constant $C(n,p,q)$ in (1.1) depends only on $n,$ $p$ and $q.$ In [17] also a weak type version of (1.1) is proved.
@@
for any sequence $\{f_{k}\}^{\infty}_{k=1}.$ Such inequalities seem to have been studied for the first time systematically by J.-L. Rubio de Francia [38]. For a thorough modern exposition of the methods in [38] we refer to [40] as well as [21].
The ideas in [38] already hinted at an intimate connection between extrapolation and bounds for vector valued extensions as in (1.2). By an extrapolation problem one understands the following. Given an operator $T$ acting on (suitable) functions on $\mathbb{R},$ we assume that for some $1<p<\infty$ it is already known or given that for all weights (that means, a.e. positive locally integrable functions) $w$ on $\mathbb{R}^{n}$ that belong to some class $C(p),$ one has the estimate $\|Tf\|_{\mathrm{L}^{p}(w)}\leq C(T,n,p,w)\|f\|_{\mathrm{L}^{p}(w)}.$ Given this information, find all $1<q<\infty$ as well as associated classes $C(q)$ of weights on $\mathbb{R}^{n},$ such that for any $w\in C(q)$ one has an estimate of the form $\|Tf\|_{\mathrm{L}^{q}(w)}\leq C(T,n,q,w)\|f\|_{\mathrm{L}^{q}(w)}.$ Rubio de Francia [37] solved completely the extrapolation problem in the case that $C(p)$ coincides with the Muckenhoupt $A_{p}$ class, that is $w\in A_{p}$ if and only if
$w]_{A_{p}}\coloneq\sup_{Q}\left(\frac{1}{|Q|}\int_{Q}w(x)\,\mathrm{dm}(x)\right)\left(\frac{1}{|Q|}\int_{Q}w(x)^{-1/(p-1)}\,\mathrm{dm}(x)\right)^{p-1}<\infty,$
where the supremum ranges again over all cubes $Q\subseteq\mathbb{R}^{n}.$ In this case, beginning with any fixed $1<p<\infty$ and an extrapolation hypothesis holding for all weights $w\in A_{p},$ the extrapolation problem is solvable for any $1<q<\infty$ and all weights $w\in A_{q}.$ The extrapolation theorem of Rubio de Francia was subsequently further refined by various authors, until a sharp quantitative version of it was proved by O. Dragičević, L. Grafakos, M. C. Pereyra and S. Petermichl [15] (see also [16]). A very thorough treatment of various forms of extrapolation with extensive historical background can be found in [10]. In fact, this method is so powerful, that it naturally yields weighted estimates for vector valued extensions, that is
@@
as explained in detail in [10].
Inequalities of the form (1.3) are a major tool for estimating operators arising naturally when decomposing biparameter operators or bicommutators in simpler, localized pieces. Such a decomposition for the so called Journeé operators was established by H. Martikainen [32] (generalizing an analogous decomposition proved earlier by T. Hytönen [24] for Calderón–Zygmound operators in the context of the solution of the $A_{2}$ problem). I. Holmes, S. Petermichl and B. Wick [23] showed that these localized pieces can be estimated in terms of a weighted product BMO space. In the following, we recall the relevant definitions.
@@
where the supremum is taken over all cubes $Q\subseteq\mathbb{R}^{n},$ and we denote $\langle b\rangle_{Q}\coloneq\frac{1}{|Q|}\int_{Q}b(x)\,\mathrm{dm}(x).$ The importance of this space is two-fold. First, it is the dual space to the (real variable) Hardy space $\mathrm{H}^{1}(\mathbb{R}^{n}).$ Second, the norm $\|b\|_{\mathrm{BMO}(\mathbb{R}^{n})}$ is the “correct” quantity controlling the boundedness of commutators $T,b]=[T,M_{b}],$ where $M_{b}$ denotes (pointwise) multiplication by $b$ (called symbol of the commutator) and $T$ is a Calderón–Zygmound operator. This was proved for the Hilbert transform by Z. Nehari [35] and in full generality by R. R. Coifmann, Rochberg and G. Weiss [8]. Moreover, the John–Nirenberg inequalities are an important intrinsic property of the space $\mathrm{BMO}(\mathbb{R}^{n}).$
B. Muckenhoupt and R. L. Wheeden [34] considered and studied the weighted BMO norm
@@
where the supremum ranges over all cubes $Q\subseteq\mathbb{R}^{n}$ and $\nu$ is a $A_{2}$ weight on $\mathbb{R}^{n}.$ A characterization of the two weighted boundedness of commutators in terms of a weighted BMO norm of the symbol was established by S. Bloom [4] for the Hilbert transform and later for arbitrary Calderón–Zygmund operators by I. Holmes, M. Lacey and B. Wick [22]. In the latter work two weighted versions of (1.4) were introduced and associated John–Nirenberg inequalities were established. These played an important role in the commutator estimates in [22].
The study of biparameter BMO spaces on product spaces $\mathbb{R}^{n}\times\mathbb{R}^{m}$ was initiated by S. Y. A. Chang [6] and R. Fefferman [18]. Here, “biparameter” refers to invariance of the considered function spaces under rescaling each coordinate variable of the domain of definition separately. The papers [6] and [18] introduced and investigated the biparameter product BMO space $\text{BMO}(\mathbb{R}\times\mathbb{R})$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{2}$ (considered as the product space $\mathbb{R}\times\mathbb{R}$) such that
@@
The aforementioned works [6] and [18] established in particular that $\text{BMO}(\mathbb{R}\times\mathbb{R})$ is the dual to the biparameter Hardy space $\mathrm{H}^{1}(\mathbb{R}\times\mathbb{R}).$ Moreover, a dyadic version of this product BMO space is the correct space for characterizing the boundedness of bicommutators $T_{1},[T_{2},b]],$ where $T_{1},T_{2}$ are Haar multipliers, as proved by [3].
A weighted version of the Chang–Fefferman product BMO space was introduced and studied by Holmes–Petermichl–Wick [23] in the context of proving two weight upper bounds for biparameter praproducts. Namely, given a biparameter dyadic grid $\bm{\mathcal{D}}$ in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and a biparameter $\bm{\mathcal{D}}$-dyadic $A_{2}$ weight $w$ on $\mathbb{R}^{n}\times\mathbb{R}^{m},$ [23] considers the dyadic Bloom type product space $\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(\nu)$ consisting of all locally integrable functions $b$ on $\mathbb{R}^{n}\times\mathbb{R}^{m}$ with the property
@@
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n}\times\mathbb{R}^{m}$ of nonzero finite measure. Section 2 contains an explanation of the notation used here, including the definition of the Haar coefficients $b_{R}^{\varepsilon},$ which can be found in expression (2.1). In [23] an $\mathrm{H}^{1}$-BMO duality type result was established in this setting, which played a crucial role in the proofs of the upper bounds there. More recently, a two-weight version of (1.5) was defined in [31] and associated John–Nirenberg inequalities were established. These played an important role in [31] for characterizing the two weight boundedness of bicommutators with Haar multipliers, extending the aforementioned result of [3] to the two weight setting.
@@ 1.1 Main results
We refer to Subsection 3.1 for a detailed explanation of the notation in Theorem 1. Theorem 1 yields readily a similar estimate for the so called Christ–Goldberg auxiliary maximal operator, as explained in Theorem 6 below.
@@
We deduce Theorem 1 from a general principle for establishing matrix weigh-ted bounds for vector valued extensions of operators acting on convex body valued functions, see Theorem 4 below. Our method for deducing such bounds is inspired from [10]: we use an analog of the Rubio de Francia extrapolation theorem for matrix weights proved in [5], coupled with a trick of interpreting vector valued extensions of operators as operators whose values are convex body valued functions. It is worth noting that using the exact same method, the recent limited range extrapolation theorem for matrix weights proved in [30] yields similar bounds as in Theorem 4 that are valid only for a limited range of exponents.
Here it is important to remark one essential difference between the extrapolation theorem proved in [5] and what we actually use. The extrapolation for matrix weights and the other methods and techniques from [5] (which also belong to the foundations of the work in [30]) concern the setting of real convex body valued functions. That is, functions taking values on the set of convex bodies in $\mathbb{R}^{d}.$ Our results and the techniques used to achieve them refer to complex valued objects, namely $\mathbb{C}^{d}$ vector valued functions and complex convex body valued functions (taking values on the set of convex bodies in $\mathbb{C}^{d}$). Most of the steps in translating the results from the real setting to the complex setting are immediate, thanks to the fact that $\mathbb{C}^{d}$ and $\mathbb{R}^{2d}$ share the same structure as topological spaces, metric spaces, measure spaces and real vector spaces. However, not all steps are covered by these considerations. In particular, the main difficulty is that it is not clear that, given a measurably parametrized family of norms over $\mathbb{C}^{d}$ (such as the ones defined by a matrix weight), there is a measurable choice of a reducing operator for such a family. This was already proved separately in [14]. Nonetheless, in Appendix A we cover the proof of Theorem A. The more obvious steps are only mentioned together with the reason why they hold exactly as in the real variable case, while we devote more detail to the steps that are not immediate.
An immediate application of Theorem 1 and Theorem 6, which is the analogue for the Christ–Goldberg auxiliary maximal operator, is the following. In [14], the authors show matrix weighted $\mathrm{L}^{p}$ bounds for Journé operators in the biparameter setting. However, the bounds in that article are only complete in the case $p=2.$ Even though that article also contains bounds for all $1<p<\infty,$ for $p\neq 2$ these only hold for the particular class of paraproduct free Journé operators. The general bounds can be obtained using extrapolation for biparameter matrix weights (see [41]). Nonetheless, one can also apply Theorem 1 and Theorem 6 as it was outlined already in [14, Section 8] to get complete bounds for all biparameter Journé operators, avoiding in this way the use of extrapolation for biparameter matrix weights.
@@
In this paper we focus on the application of Theorem 1 for setting up the foundations of a theory of two matrix weighted product BMO. Namely, let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in the product space $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We define
@@
where the supremum ranges over all Lebesgue-measurable subsets $\Omega$ of $\mathbb{R}^{n+m}$ of nonzero finite measure. This definition is an extension of one of the equivalent definitions for the space of two matrix weighted one-parameter BMO, whose study was initiated in [25], [28] and culminated in [27]. Moreover, for every sequence $\Phi=\{\Phi_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C}),$ we define
@@
which is the direct biparameter analog of the one parameter two matrix weighted $\mathrm{H}^{1}$ norm from [28]. In this context, we prove the following result.
@@
is well-defined and bounded with $\|\ell_{B}\|\sim\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)},$ where the implicit constants depend only on $n,m,p,d$ and $V]_{A_{p},\bm{\mathcal{D}}}.$ Conversely, for every bounded linear functional $\ell$ on $\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ there is $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)$ with $\ell=\ell_{B}.$
We note that our proof of Theorem 2 trivially works also in the one-parameter setting, thus answering to the positive the question posed in [28] about the extension of the one parameter two matrix weighted $\mathrm{H}^{1}$-BMO duality proved there from $p=2$ to arbitrary exponents $1<p<\infty.$
The duality result of Theorem 2 coupled with Theorem 1 yields readily two-matrix weighted bounds for biparameter paraproducts, see Proposition 10 and Proposition 13 below. These in turn yield some two matrix weighted upper bounds for bicommutators, see Subsection 4.3 below.
The remainder of this article is structured as follows. In Section 2, we review the objects that we work with and their relevant properties for our results. Section 3 is focused on vector valued estimates. There we state the relevant result of [5] in the complex setting, we outline very briefly its proof (deferring the more technical details to the Appendix), and then we use it to show vector valued estimates for general families of extrapolation pairs. This general result is followed by vector valued estimates for matrix weighted maximal operators. Namely, we show such estimates both for the Christ–Goldberg maximal operator and for the Christ–Goldberg auxiliary maximal operator. In Section 4 we develop a two-matrix weighted $\mathrm{BMO}$ theory. First, we define a two-matrix weighted matrix $\mathrm{BMO}$ space, and then we show that it is dual to a two-matrix weighted matrix $\mathrm{H}^{1}$ space already considered in the literature. Next, we use the previously developed vector valued estimates to prove norm bounds for biparameter matrix paraproducts, relating them to the $\mathrm{BMO}$ norm of their symbol. We conclude that section with a sketch on how our results can be also applied to obtaining bounds for bicommutators with Haar multipliers and multiplication with a matrix symbol. Lastly, in Appendix A we include the more technical details on how to obtain the complex version of Bownik and Cruz-Uribe extrapolation for matrix weights found in [5].